Grade 12IIT JEE Entrance Exam

For what value of x is x^2-3x-18 positive?

Rickey
11 Years agoGrade 12

1 Answer

Vikas TU
ApprovedApproved Tutor Answer11 Years ago
You have gotten this far: 
+++++(-3)----------(6)++++++ 

Consider the intervals (-∞, -3),(-3,6),(6, ∞) 
Choose 'any' one point from each of the intervals. 

(-∞, -3) : choose x=-4 
x^2-3x-18 where x=-4 = 10 > 0 (positive) 

(-3,6): choose x=0 
x^2-3x-18 where x=0 = -18

(6, ∞) : choose x=7 
x^2-3x-18 where x=7 = 10 > 0 (positive) 

The intervals where x is positive are: 
(-∞, -3) and (6, ∞) 

(-∞, -3) U (6, ∞) is the answer 
or 
x 6